If the letters of the word MASTER are permuted in all possible ways and the words thus
formed are arranged in the dictionary order, then find the rank of the word '' REMAST ''.

Grade:upto college level

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
The alphabetical order of the letters of the given word is A, E, M, R, S, T
The number of words begin with A is 5! = 120
The number of words begin with E is 5! = 120
The number of words begin with M is 5! =120
The number of words begin with RA is 4! = 24
The number of words begin with REA is 3! = 6
The next word is REMAST .
∴Rank of the word REMAST - 3 (120) + 24 + 6 + 1 = 360 + 31 = 391
Thanks and Regards
Shaik Aasif
askIITians faculty